# Quantum cellular automata and categorical dualities of spin chains

### v1

@@ Abstract.
Dualities play a central role in the study of quantum spin chains, providing insight into the structure of quantum phase diagrams and phase transitions. In this work we study categorical dualities, which are defined as bounded-spread isomorphisms between algebras of symmetry-respecting local operators on a spin chain. We consider generalized global symmetries that correspond to unitary fusion categories, which are represented by matrix-product operator algebras. A fundamental question about dualities is whether they can be extended to quantum cellular automata on the larger algebra generated by all local operators that respect the unit matrix-product operator. For conventional global symmetries, which are on-site representations of finite groups, this larger algebra is simply the tensor product of algebras associated to individual spins in the chain. We present a solution to the extension problem using the machinery of Doplicher-Haag-Roberts bimodules. Our solution provides a crisp categorical criterion for when an extension of a duality exists. We show that the set of possible extensions form a torsor over the invertible objects in the relevant symmetry category. Our approach recovers existing results from the literature when applied to quantum cellular automata that respect a conventional global symmetry.
@@ 1. Introduction
Thus far, most work has focused on QCAs embodied by bounded-spread automorphisms on standard quasi-local operator algebras of spin systems $A=\otimes_{\mathbbm{Z}^{n}}M_{d}(\mathbbm{C})$, formed by taking the tensor product of finite qudit degrees of freedom on the sites of a lattice. This has led to a full classification of QCAs in one spatial dimension [GNVW09] and partial classifications in higher spatial dimensions [FH20, Haa21, Haa22]. There is a closely related problem of understanding dualities, which in this context are mathematically described by bounded-spread isomorphisms between subalgebras of the usual quasi-local algebra that interchange quantum phases of matter. Classification problems in this direction remain largely open (see  [AMF16, TW19, CCH+21, KOZ22, LDOV23, Jon24, JL24, BDSNY24] for recent progress), and it is these dualities in one spatial dimension that form the primary focus of this work.
@@
Dualities arise naturally in the presence of a global symmetry $\mathcal{C}\curvearrowright A$, which may correspond to a group, (weak) Hopf algebra, or fusion category. If a pair of Hamiltonians has symmetric local terms, then these lie in the algebra $A^{\mathcal{C}}\subseteq A$ of symmetric operators. In this work, we use the term duality to formally denote a bounded-spread isomorphism between subalegbras of symmetric operators $A^{\mathcal{C}}$ under some global symmetry which exchange Hamiltonians but may not extend to bounded-spread automorphisms of the whole quasi-local algebra $A$, also known as QCA. The motivating example for this class of dualities is due to Kramers and Wannier, originally introduced in the context of classical statistical mechanics  [KW41]. The bounded-spread automorphism implementing KW duality acts on the symmetric subalgebra of operators on a chain of qubits that commute with the global $\mathbb{Z}_{2}$ symmetry represented by $\prod_{i}\sigma^{x}_{i}$. The action of the Kramers-Wannier (KW) duality can be specified on a set of generators of this $\mathbb{Z}_{2}$-symmetric operator algebra via the map $\sigma^{x}_{i}\mapsto\sigma^{z}_{i-1}\sigma^{z}_{i}$ and $\sigma^{z}_{i}\sigma^{z}_{i+1}\mapsto\sigma^{x}_{i}$.
@@
In recent years, there has been significant interest in extending results from the setting of conventional global symmetries, which are represented by tensor products of single-site unitary operators, to categorical global symmetries, which are represented by matrix product operators (MPOs) [FFRS04, BMW+17, VBW+18, AMF16, TW19]. The resulting algebra generated by symmetric local operators is called a fusion spin chain. Recently, the idea of KW duality has been generalized in this direction to categorical symmetries, where matrix-product operators have been found to implement dualities based on any invertible bimodule categories between fusion categories [AMF16, WBV17, AFM20, LDOV23]. This suggests a deep connection between categorical dualities and the algebraic structure of fusion categories.
@@
A key feature of non-trivial dualities following the KW paradigm is that bounded-spread isomorphisms of the symmetric operators should not be locally extendable to the whole spin chain (or in the case of non-unital MPOs, to the local edge-restricted sector). We call such an extension a spatial implementation. Spatial implementations are themselves ordinary 1+1D QCAs in the case of group or Hopf algebra symmetries. More generally, spatial implementations are “edge-restricted” QCA. These extensions lead naturally to the following question, whose answer sheds light on the relationship between QCA and categorical dualities:
@@
When do bounded-spread isomorphisms of symmetric local operator algebras (i.e. categorical dualities) extend to QCA defined on the full (or edge-restricted) local operator algebras, and how are the possible extensions characterized?
To answer this question, we first establish a precise reformulation in terms of abstract fusion spin chains $A(\mathcal{E},X)$, built from a fusion category $\mathcal{E}$ and an object $X\in\mathcal{E}$, by defining local algebras as endomorphism algebras of tensor powers of $X$ (see Section 3). A fusion category $\mathcal{C}$ acting on a spin chain via MPOs is specified by the data of an indecomposable right $\mathcal{C}$-module category $\mathcal{M}$, and an object $X\in\mathcal{C}^{*}_{\mathcal{M}}$. Then, the symmetric operator algebra is isomorphic to the fusion spin chain $A(\mathcal{C}^{*}_{\mathcal{M}},X)$. If we view $\mathcal{M}$ as a left $\mathcal{C}^{*}_{\mathcal{M}}$-module category, then rewriting $\mathcal{E}:=\mathcal{D}^{*}_{\mathcal{M}}$ we can reformulate the data of a categorical symmetry as an abstract fusion spin chain $A(\mathcal{C},X)$ together with an indecomposable left $\mathcal{C}$-module category $\mathcal{M}$. The inclusion $A(\mathcal{E},X)\hookrightarrow A(\mathcal{E},X)_{\mathcal{M}}$ (see Definition 3.3) corresponds to the inclusion of symmetric operators into all (edge-restricted) local operators, which we call a spatial realization of the chain $A(\mathcal{E},X)$. Question 1.1 can be given a precise mathematical formulation as follows:
@@
Given a bounded-spread isomorphism $\alpha:A(\mathcal{C},X)\rightarrow A(\mathcal{D},Y)$ when does it extend to a QCA between spatial realizations $A(\mathcal{C},X)_{\mathcal{M}}\rightarrow A(\mathcal{D},Y)_{\mathcal{N}}$, and how are these extensions characterized?
@@
For the Kramers-Wannier example described above, the fusion category is $\mathcal{C}=\mathcal{D}=\text{Rep}(\mathbbm{Z}_{2})$, $X$ is the left regular representation of $\mathbbm{Z}_{2}$, and the DHR category of the symmetric spin chain is the toric code modular tensor category $\mathcal{Z}(\text{Hilb}_{f.d.})$. The module category in question is the standard fiber functor, which corresponds to the electric Lagrangian $L=1\oplus e$. The induced action of the Kramer-Wannier automorphism $\alpha$ is the $e\leftrightarrow m$ swap symmetry, which does not preserve $L$. By the above theorem, $\alpha$ admits no bounded-spread extension to the whole quasi-local algebra $A=\otimes_{\mathbbm{Z}}M_{2}(\mathbbm{C})$.
Finally, we can use our formalism to make contact with the literature concerning symmetric QCA (sQCA). In this setting, we consider a global on-site finite group $G$ acting by unitaries. We have the following (c.f. [GSSC20]) as a consequence of our machinery (see Theorem 6.3):
Theorem 1.4.
There is a homomorphism $\pi:\text{sQCA}\rightarrow H^{2}(G,U(1))$.
Our results on the extension of dualities to QCAs point to a number of directions for future work. First, what is the full classification of bounded-spread isomorphisms on local algebras and similarly what is the classification of edge-restricted QCAs in one spatial dimension? Second, can our results be applied to characterize symmetries of topological orders in two dimensions via their action on the boundary? Third, can techniques from the theory of operator algebras, subfactors, and fusion categories be applied to classify dualities and QCAs in higher spatial dimensions?
The remainder of this work is presented as follows. In section 2 we introduce background material about categorical dualities of quantum spin chains. In section 3 we describe dualities in the setting of fusion spin chains. In section 4 we describe categorical inclusions defined by module categories over algebra objects. In section 5 we describe how DHR bimodules can be applied to solve the extension problem for dualities of quantum spin chains. In section 6 we present examples that generalize the KW duality.
@@ 2. Dualities for spin chains with categorical symmetry
$\sum_{\{i\}\{j\}}\operatorname{tr}(A^{i_{1}j_{1}}A^{i_{2}j_{2}}...\noindent A^{i_{n}j_{n}})\ket{\{i\}}\bra{\{j\}}$ | (2.6)
where
@@
where we treat $A^{ij}_{\alpha\beta}$ as a matrix element of $A^{ij}$ and take a matrix product over the $\alpha,\beta$ indices of the tensors $A^{ij}_{\alpha\beta}$. From the data in Definition 2.4, one can construct for each simple object in $\mathcal{C}$ a matrix product operator, satisfying a family of categorical relations [SWB+14, BMW+17]. It is in this sense that a fusion category is actually acting on the spin chain.
@@
$H=\sum_{i\in\mathbbm{Z}}J\sigma^{z}_{i}\sigma^{z}_{i+1}+h\sigma^{x}_{i},$
@@ 3. Abstract fusion spin chains
If we change our perspective and denote $\mathcal{E}:=\mathcal{C}^{*}_{\mathcal{M}}$, then we can view $\mathcal{M}$ as a left indecomposable $\mathcal{E}$-module category, and thus spatial realizations of the fusion spin chain $A(\mathcal{E},X)$, parameterized by a left module category $\mathcal{M}$, correspond precisely the algebra symmetry operators inside $A^{\circ}$ for an action of the dual category $\mathcal{C}^{*}_{\mathcal{M}}\curvearrowright A$. This leads us to reformulate Question 2.10 as follows:
@@ 5. Solving the extension problem with DHR bimodules
If $X$ is a strong, self-dual tensor generator of the unitary fusion category $\mathcal{C}$, then $\text{DHR}({\mathfrak{A}}(\mathcal{C},X)\cong\mathcal{Z}(\mathcal{C})$ as braided C*-tensor categories.
@@
    If $\text{DHR}(\alpha)(L)$ is not isomorphic to $Z(K)$, then $\alpha$ has no extension to a bounded-spread isomorphism $\widetilde{\alpha}:A(\mathcal{C}_{L},X)\rightarrow A(\mathcal{D}_{K},Y)$
@@
Proof of Theorem 1.3. The Theorem 1.3 stated in the introduction follows as a special case of the above, and we believe it provides a satisfactory answer to Question 2.10. Indeed, consider the case where $L\in\mathcal{Z}(\mathcal{C})$ and $K\in\mathcal{Z}(\mathcal{D})$ are $Z(\mathcal{M})$ and $Z(\mathcal{N})$ associated to the $\mathcal{C}$ and $\mathcal{D}$ module categories, respectively. Then we have identifications of the spatial realizations $A(\mathcal{C},X)_{\mathcal{M}}\cong A(\mathcal{C}_{Z(\mathcal{M})},X)$, whence the first part of the result follows from the previous theorem. That $\text{Aut}(Z(\mathcal{M}))\cong(\mathcal{C}^{*}_{\mathcal{M}})^{\times}$ follows, for example, from [BJ22]. ∎

### v2

@@ Abstract.
Dualities play a central role in the study of quantum spin chains, providing insight into the structure of quantum phase diagrams and phase transitions. In this work, we study categorical dualities, which are defined as bounded-spread isomorphisms between algebras of symmetry-respecting local operators on a spin chain. We consider generalized global symmetries that correspond to unitary fusion categories, which are represented by matrix-product operator algebras. A fundamental question about dualities is whether they can be extended to quantum cellular automata on the larger algebra generated by all local operators in the the unit matrix-product operator sector. For on-site representations of Hopf algebra symmetries, this larger algebra is the usual tensor product quasi-local algebra. We present a solution to the extension problem using the machinery of Doplicher-Haag-Roberts bimodules. Our solution provides a crisp categorical criterion for when an extension of a duality exists. We show that the set of possible extensions form a torsor over the invertible objects in the relevant symmetry category. As a corollary, we obtain a classification result concerning dualities in the group case.
@@ 1. Introduction
Thus far, most work has focused on QCAs embodied by bounded-spread automorphisms on the standard quasi-local operator algebra of a spin system, $A=\otimes_{\mathbbm{Z}^{n}}M_{d}(\mathbbm{C})$, formed by taking the tensor product of finite qudit degrees of freedom on the sites of a lattice. This has led to a full classification of QCAs in one spatial dimension up to finite-depth circuits  [GNVW09] and partial classifications in higher spatial dimensions [FH20, Haa21, Haa22]. There is a closely related problem of understanding dualities, which in this context are mathematically described by bounded-spread isomorphisms between subalgebras of the usual quasi-local algebra that interchange quantum phases of matter. Questions in this direction have attracted much recent interest [AMF16, TW19, CCH+21, KOZ22, LDOV23, Jon24, JL24, BDSNY24, MLC24].
@@
Dualities can arise naturally in the presence of a global symmetry $\mathcal{C}\curvearrowright A$, which may correspond to an action of a group, a (weak) Hopf algebra, or a fusion category. If a pair of Hamiltonians has symmetric local terms, then these lie in the algebra $A^{\mathcal{C}}\subseteq A$ of symmetric operators. In this work, we use the term duality to formally denote a bounded-spread isomorphism between subalgebras of symmetric operators $A^{\mathcal{C}}$ under some global symmetry which exchange Hamiltonians but may not extend to bounded-spread automorphisms of the whole quasi-local algebra $A$ (QCA). The motivating example for this class of dualities is due to Kramers and Wannier, originally introduced in the context of classical statistical mechanics  [KW41]. The bounded-spread automorphism implementing KW duality acts on the symmetric subalgebra of operators on a chain of qubits that commute with the global $\mathbb{Z}_{2}$ symmetry represented by $\prod_{i}\sigma^{x}_{i}$. The action of the Kramers-Wannier (KW) duality can be specified on a set of generators of this $\mathbb{Z}_{2}$-symmetric operator algebra via the map $\sigma^{x}_{i}\mapsto\sigma^{z}_{i-1}\sigma^{z}_{i}$ and $\sigma^{z}_{i}\sigma^{z}_{i+1}\mapsto\sigma^{x}_{i}$.
@@
In recent years, there has been significant interest in extending results from the setting of conventional global symmetries, which are represented by tensor products of single-site unitary operators, to categorical global symmetries, which are represented by matrix product operators (MPOs) [FFRS04, BMW+17, VBW+18, AMF16, TW19]. In this work, we study this flavor of duality in one spatial dimension. The resulting algebra generated by symmetric local operators (preserving the unit MPO sector) is called a fusion spin chain. Recently, the idea of KW duality has been generalized in this direction to categorical symmetries, where matrix-product operators have been found to implement dualities based on any invertible bimodule categories between fusion categories [AMF16, WBV17, AFM20, LDOV23]. This suggests a deep connection between categorical dualities and the algebraic structure of fusion categories.
@@
A key feature of non-trivial dualities following the KW paradigm is that bounded-spread isomorphisms of the symmetric operators should not be locally extendable to the whole spin chain (or in the case of non-unital MPOs, to the local edge-restricted sector). We call such an extension a spatial implementation. Spatial implementations are themselves ordinary 1+1D QCAs in the case of group or Hopf algebra symmetries. More generally, spatial implementations are edge-restricted QCA. These extensions lead naturally to the following question, whose answer sheds light on the relationship between QCA and categorical dualities:
@@
When do bounded-spread isomorphisms between symmetric local operator algebras (i.e. categorical dualities) extend to QCA defined on the full (or edge-restricted) local operator algebras, and how are the possible extensions characterized?
To answer this question, we first establish a precise reformulation in terms of abstract fusion spin chains $A(\mathcal{E},X)$, built from a fusion category $\mathcal{E}$ and an object $X\in\mathcal{E}$, by defining local algebras as endomorphism algebras of tensor powers of $X$ (see Section 3). A fusion category $\mathcal{C}$ acting on a spin chain via MPOs is specified by the data of an indecomposable right $\mathcal{C}$-module category $\mathcal{M}$, and an object $X\in\mathcal{C}^{*}_{\mathcal{M}}$. Then, the symmetric operator algebra is isomorphic to the fusion spin chain $A(\mathcal{C}^{*}_{\mathcal{M}},X)$. If we view $\mathcal{M}$ as a left $\mathcal{C}^{*}_{\mathcal{M}}$-module category, then rewriting $\mathcal{E}:=\mathcal{C}^{*}_{\mathcal{M}}$ we can reformulate the data of a categorical symmetry as an abstract fusion spin chain $A(\mathcal{C},X)$ together with an indecomposable left $\mathcal{C}$-module category $\mathcal{M}$. The inclusion $A(\mathcal{E},X)\hookrightarrow A(\mathcal{E},X)_{\mathcal{M}}$ (see Definition 3.3) corresponds to the inclusion of symmetric operators into all (edge-restricted) local operators, which we call a spatial realization of the chain $A(\mathcal{E},X)$. Question 1.1 can be given a precise mathematical formulation as follows:
@@
Given a bounded-spread isomorphism $\alpha:A(\mathcal{C},X)\rightarrow A(\mathcal{D},Y)$ between abstract fusion spin chains, when does it extend to a QCA between spatial realizations $A(\mathcal{C},X)_{\mathcal{M}}\rightarrow A(\mathcal{D},Y)_{\mathcal{N}}$, and how are these extensions characterized?
@@
For the Kramers-Wannier example described above, the fusion category is $\mathcal{C}=\mathcal{D}=\text{Rep}(\mathbbm{Z}_{2})$, $X$ is the left regular representation of $\mathbbm{Z}_{2}$, and the DHR category of the symmetric spin chain is the toric code modular tensor category $\mathcal{Z}(\text{Hilb}_{f.d.}(\mathbbm{Z}_{2}))$. The module category in question is the standard fiber functor, which corresponds to the electric Lagrangian $L=1\oplus e$. The induced action of the Kramer-Wannier automorphism $\alpha$ is the $e\leftrightarrow m$ swap symmetry, which does not preserve $L$. By the above theorem, $\alpha$ admits no QCA extension to the whole quasi-local algebra $A=\otimes_{\mathbbm{Z}}M_{2}(\mathbbm{C})$.
We can also use our formalism to make contact with the literature concerning symmetric QCA (sQCA). In this setting, we consider a global, on-site finite group $G$ acting faithfully by unitaries in a self-dual representation. Then the symmetric operators $A^{G}$ have DHR category equivalent to the Drinfeld center $\mathcal{Z}(\text{Rep}(G))$. We can immediately recover an $H^{2}(G,U(1))$ index (c.f. [GSSC20]) as a consequence of our framework (see Theorem 6.3).
[…]
Corollary 1.4.
[…]
In other words, for an onsite group symmetry the DHR action of a duality on the quantum double completely classifies dualities up to finite-depth circuits that are globally symmetric, and have trivial $H^{2}(G,U(1))$ index. In general, the difference between globally symmetric finite-depth circuits with trivial $H^{2}(G,U(1))$ index and locally symmetric finite-depth circuits is unclear. Here, a locally symmetric finite-depth circuit must have some decomposition into local quantum gates that individually commute with the symmetry, whereas a globally symmetric finite-depth circuit only needs to commute with the symmetry globally. Some progress in this direction has been made in [GSSC20], utilizing a generalized GNVW index. Unfortunatey these results do not map precisely onto our framework. However, for abelian groups in the regular representation, it was recently shown that globally symmetric finite depth circuits with trivial $H^{2}(G,U(1))$ index are always locally symmetric [MLC24, Theorem 1]. Corollary 1.4 combined with this result then yields an alternate version of the main result in [MLC24], which uses a different version of the index.
Our results point to a number of directions for future work. First, what is the full classification of bounded-spread isomorphisms on local algebras for fusion spin chains? Using our results, we expect to be able to reduce this problem to the classification of edge-restricted QCAs in one spatial dimension, which is an open problem. Second, can our results be applied to characterize symmetries of topological orders in two dimensions via their action by dualities on the holographic boundary (for example, in the sense of [JNPW23])? Third, can techniques from the theory of operator algebras, subfactors, and fusion categories be applied to classify dualities and QCAs in higher spatial dimensions?
The remainder of this work is presented as follows. In Section 2 we introduce background material about categorical dualities of quantum spin chains. In Section 3 we describe dualities in the setting of fusion spin chains. In Section 4 we describe categorical inclusions defined by module categories over algebra objects. In Section 5 we describe how DHR bimodules can be applied to solve the extension problem for dualities of quantum spin chains. In Section 6 we present examples that generalize the KW duality, and give our applications.
@@ 2. Dualities for spin chains with categorical symmetry
$\sum_{\{i\}\{j\}}\operatorname{tr}(M^{i_{1}j_{1}}M^{i_{2}j_{2}}...\noindent M^{i_{n}j_{n}})\ket{\{i\}}\bra{\{j\}}$ | (2.6)
where $\{i\}=\{i_{1},i_{2},\dots,i_{n}\}$ and similarly for $\{j\}$, and
@@
where we treat $M^{ij}_{\alpha\beta}$ as a matrix element of $M^{ij}$ and take a matrix product over the $\alpha,\beta$ indices of the tensors $M^{ij}_{\alpha\beta}$. From the data in Definition 2.4, one can construct for each (simple) object in $\mathcal{C}$ a matrix product operator, satisfying a family of categorical relations [SWB+14, BMW+17]. It is in this sense that a fusion category is actually acting on the spin chain. Alternatively, one can build from an MPO a locality preserving quantum channel, which is defined on local operators via a suitable normalization of the following picture
where $M^{*}$ denotes the complex conjugate of $M$ with the $i$ and $j$ indices swapped. The categorical relations imply that these quantum channels implement a normalized version of the fusion rules of the category.
@@
$H_{J,h}=\sum_{i\in\mathbbm{Z}}J\sigma^{z}_{i}\sigma^{z}_{i+1}+h\sigma^{x}_{i},$
@@ 3. Abstract fusion spin chains
If we change our perspective and denote $\mathcal{E}:=\mathcal{C}^{*}_{\mathcal{M}}$, then we can view $\mathcal{M}$ as a left indecomposable $\mathcal{E}$-module category, and thus spatial realizations of the fusion spin chain $A(\mathcal{E},X)$, parameterized by a left module category $\mathcal{M}$, correspond precisely the algebra symmetry operators inside $A^{\circ}$ for an action of the dual category $\mathcal{C}^{*}_{\mathcal{M}}\curvearrowright A$. This leads us to reformulate Question 2.11 as follows:
@@ 5. Solving the extension problem with DHR bimodules
If $X$ is a strong, self-dual tensor generator of the unitary fusion category $\mathcal{C}$, then $\text{DHR}({\mathfrak{A}}(\mathcal{C},X))\cong\mathcal{Z}(\mathcal{C})$ as braided C*-tensor categories.
The above theorem provides an MPO picture for DHR bimodules, in the case of a spatial realization. First, for an object $Z\in\mathcal{Z}(\mathcal{C})$, we consider defect tensors $a$ as in the following picture, where the horizontal string is labeled by an object $Z$ in the Drinfeld center, and $M$ is the associated MPO for the object $Z\in\mathcal{C}$.
See [BMW+17, WBV17] for a description of how to decompose the space of defect tensors for MPOs corresponding to a category $\mathcal{C}$ into simple objects of the center $\mathcal{Z}(\mathcal{C})$.
Next, we can represent the half-braiding associated to $Z$ with an MPO $h$, and in this way extend the action of the fusion ring by channels on local operators to the defect tensors via the following picture.
The defect sectors are decomposed into irreducible representations under this action, which is similar to Ocneanu’s tube algebra [Ocn88].
Finally, we consider the vector space of all such defect tensors invariant under the $\mathcal{C}$-symmetry. This is a bimodule under the action of symmetric local operators $L_{1},L_{2},\dots$ and $R_{1},R_{2},\dots$ as depicted in the following picture.
By the above theorem, it turns out that these constitute all the abstractly defined DHR bimodules in the symmetric case.
@@
    If $\text{DHR}(\alpha)(L)$ is not isomorphic to $K$, then $\alpha$ has no extension to a bounded-spread isomorphism $\widetilde{\alpha}:A(\mathcal{C}_{L},X)\rightarrow A(\mathcal{D}_{K},Y)$
@@
Proof of Theorem 1.3. The Theorem 1.3 stated in the introduction follows as a special case of the above, and we believe it provides a satisfactory answer to Question 2.11. Indeed, consider the case where $L\in\mathcal{Z}(\mathcal{C})$ and $K\in\mathcal{Z}(\mathcal{D})$ are $Z(\mathcal{M})$ and $Z(\mathcal{N})$ associated to the $\mathcal{C}$ and $\mathcal{D}$ module categories, respectively. Then we have identifications of the spatial realizations $A(\mathcal{C},X)_{\mathcal{M}}\cong A(\mathcal{C}_{Z(\mathcal{M})},X)$, whence the first part of the result follows from the previous theorem. That $\text{Aut}(Z(\mathcal{M}))\cong(\mathcal{C}^{*}_{\mathcal{M}})^{\times}$ follows, for example, from [BJ22]. ∎